# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v2

@@ 6 Conclusion
This paper advances the field of hyperbolic LVMs by augmenting the hyperbolic manifold with a Riemannian pullback metric that combines the hyperbolic geometry with the geometry of the data distribution. By minimizing the curve energy on the pullback metric, we computed geodesics that adhere to the data manifold in the hyperbolic latent space. To do so, we addressed the limitations of current auto-differentiation techniques by providing analytical solutions for the hyperbolic SE kernel derivatives. Via multiple experiments, we demonstrated that pullback geodesics outperform standard hyperbolic geodesics.
It is worth noting that the benefits of pullback geodesics are most evident when the data exhibits smooth transitions. They become less effective when the data is inherently comprised of distinct clusters, as the pullback geodesics cannot cross high-energy regions, i.e., the data manifold boundaries. This paper focused on 2D and 3D hyperbolic latent spaces as they are easier to visualize and computationally less expensive than higher-dimensional latent spaces, while benefiting from the increased capacity of hyperbolic spaces to embed hierarchical data. Future work will investigate extensions to higher-dimensional hyperbolic latent spaces, which require to deal with more complex expressions for hyperbolic kernels. Additionally, the need for computing manual kernel derivatives is unsatisfactory. Future work will explore alternative autodifferentiation techniques built on KeOps (Feydy et al., 2020) to overcome this practical issue. Finally, the approximation of the 2D hyperbolic SE kernel is computationally expensive. Performance could potentially be increased by exploring different sampling strategies, e.g., by sampling from a Rayleigh distribution rather than from a Gaussian.
Acknowledgements
This work was partially supported by the Carl Zeiss Foundation under the project JuBot, by the European Union’s Horizon Europe Framework Programme under grant agreement No 101070596 (euROBIN), and by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ Appendix B Hyperbolic Kernel Derivatives
This section provides the complete derivations of the hyperbolic SE kernel derivatives that are necessary to compute the pullback metric tensor and to optimize geodesics on it. Eq. (22) defines the pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ based on the Jacobian mean $\bm{\mu}_{J}$ (19) and covariance matrix $\bm{\Sigma}_{J}$ (20), which in turn require the kernel derivatives $\partial k(\bm{x}^{*},\bm{X})=\frac{\partial}{\partial\bm{x}^{*}}k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N}$ and $\partial^{2}k(\bm{x}^{*},\bm{x}^{*})=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}\in\mathbb{R}^{D_{x}\times D_{x}}$. Additionally, minimizing the curve energy (23), requires the derivative of the pullback metric tensor, i.e., the derivative of $\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J}$, which requires the kernel derivative $\dfrac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})$ and $\dfrac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})$. Specifically, we have
@@
where we have $\bm{\mu}_{J}\in\mathbb{R}^{D_{y}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}\in\mathbb{R}^{D_{y}\times D_{x}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N\times D_{x}}$, and ×$\times_{n}$ denotes the $n$-th mode tensor product (Kolda & Bader, 2009a). Next, we provide the details of the derivation of the kernel derivatives for $\mathbb{H}^{2}_{\mathcal{L}}$ and $\mathbb{H}^{3}_{\mathcal{L}}$.
@@
As discussed in Sec. 4.1, the derivatives of the 2D hyperbolic SE kernel (6) are entirely determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$. Eqs (25), (26), (27), and (28) show that these derivatives in turn are defined by the first and second derivatives of $\phi_{l}(\bm{x}_{\mathcal{P}})=e^{(1+2s_{l}i)\langle\bm{x}_{\mathcal{P}},\bm{b}_{l}\rangle}$, which are computed as
$\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}})$ $=(1+2s_{l}i)\,\phi_{l}(\bm{x}_{\mathcal{P}})\left[\dfrac{\partial}{\partial\bm{x}}\bm{x}_{\mathcal{P}}\right]^{\mathsf{T}}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right],$ (46)
$\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}})$ $=(1+2s_{l}i)\Bigg{(}\left[\dfrac{\partial}{\partial\bm{x}}\bm{x}_{\mathcal{P}}\right]^{\mathsf{T}}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right]\left[\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}})\right]^{\mathsf{T}}$ (47)
@@
$\left(\dfrac{\partial}{\partial\bm{x}}\bm{x}_{\mathcal{P}}\right)_{i,j}$ =$=\begin{cases}-\frac{x_{i+1}}{(1+x_{0})^{2}}&j=0\\ \quad\frac{1}{1+x_{0}}&j=i+1\\ \quad 0&\text{otherwise}\end{cases}\qquad,\qquad\left(\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\bm{x}_{\mathcal{P}}\right)_{i,j,k}=\begin{cases}\quad\frac{2x_{i+1}}{(1+x_{0})^{3}}&j=k=0\\ -\frac{1}{(1+x_{0})^{2}}&j=0\text{ and }k=i+1\\ -\frac{1}{(1+x_{0})^{2}}&k=0\text{ and }j=i+1\\ \quad 0&\text{otherwise}\end{cases}\,,$ (51)
@@
This section provides the derivations of the 3D hyperbolic SE kernel derivatives. Sec. 4.2 introduced the helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$ and $s=\sqrt{u^{2}-1}$ with derivatives $\frac{\partial u}{\partial\bm{x}}=\bm{G}^{\mathcal{L}}\bm{z}$, $\frac{\text{d}\rho}{\text{d}u}=\frac{-1}{s}$, and $\frac{\text{d}s}{\text{d}u}=\frac{u}{s}$. Using these expressions, the first and second derivatives of the 3D hyperbolic SE kernel $k(\bm{x},\bm{y})=\frac{\rho}{s}\,e^{-\frac{\rho^{2}}{\nu}}$ are given as,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}\rho\right]\dfrac{1}{s}e^{-\frac{\rho^{2}}{\nu}}+\rho\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}s^{-1}\right]e^{-\frac{\rho^{2}}{\nu}}+\dfrac{\rho}{s}\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\nu}}\right],$ (52)
$=-\dfrac{1}{s^{2}}\bm{G}^{\mathcal{L}}\bm{z}e^{-\frac{\rho^{2}}{\nu}}-\dfrac{\rho u}{s^{3}}\bm{G}^{\mathcal{L}}\bm{z}e^{-\frac{\rho^{2}}{\nu}}+\dfrac{\rho}{s^{2}}\bm{G}^{\mathcal{L}}\bm{z}\dfrac{2\rho}{\nu}e^{-\frac{\rho^{2}}{\nu}},$ (53)
@@
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\bm{Gz}\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\nu}}+\bm{Gz}g(u)\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\nu}}\right]+g(u)e^{-\frac{\rho^{2}}{\nu}}\left[\dfrac{\partial}{\partial\bm{z}}\bm{Gz}\right],$ (55)
$=\left[\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\nu}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)\dfrac{-1}{s}\dfrac{-2\rho}{\nu}e^{-\frac{\rho^{2}}{\nu}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)e^{-\frac{\rho^{2}}{\nu}}\bm{G}^{\mathcal{L}},$ (56)
@@
$\dfrac{\partial^{2}}{\partial\bm{x}\partial\bm{x}}k(\bm{x},\bm{z})$ $=h(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{z}^{\mathsf{T}}\bm{G}^{\mathcal{L}}e^{-\frac{\rho^{2}}{\nu}}\,.$ (58)
For the third-order derivatives, we consider each column $\bm{K}_{j}\in\mathbb{R}^{4\times 1}$ and row $\bm{K}_{i}\in\mathbb{R}^{1\times 4}$ of $\bm{K}=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})$ individually, and we proceed as follows,
@@
$\dfrac{\partial}{\partial\bm{x}}\bm{K}_{i}$ $=\left(q(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}z_{i}\,+h(u)\bm{G}^{\mathcal{L}}z_{i}\,+h(u)\bm{Gz}\bm{e}_{i}^{\mathsf{T}}\right)\Delta_{1i}e^{-\frac{\rho^{2}}{\nu}}\,.$ (63)
By stacking the row derivatives $\frac{\partial}{\partial\bm{x}}\bm{K}_{i}\in\mathbb{R}^{1\times 4\times 4}$ along the first dimension and the column derivatives $\frac{\partial}{\partial\bm{z}}\bm{K}_{j}\in\mathbb{R}^{4\times 1\times 4}$ along the second, we get the full third-order kernel derivatives $\frac{\partial}{\partial\bm{x}}\bm{K},\frac{\partial}{\partial\bm{z}}\bm{K}\in\mathbb{R}^{4\times 4\times 4}$. As discussed in Sec. 4.2, automatic differentiation tools such as PyTorch cannot evaluate the helper functions $g(u)$, $h(u)$, and $q(u)$ as $u$ approaches $-1$ from below. To address this issue, we replace the general function definitions with their analytical limits, when $u$ gets too close to $-1$. This is equivalent to the hyperbolic distance between the two kernel inputs falling below a predefined threshold value $\rho=\text{arccosh}(-u)=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})\leq 1\mathrm{e}{-4}$. The analytical limit expressions for the helper functions are given as

### v3

@@ 6 Conclusion
This paper advances the field of hyperbolic LVMs by augmenting the hyperbolic manifold with a Riemannian pullback metric, combining in a principled manner the hyperbolic geometry with the geometry of the data distribution. By minimizing the curve energy on the pullback metric, we computed geodesics that adhere to the data manifold in the hyperbolic latent space. To do so, we addressed the limitations of current auto-differentiation techniques by providing analytical solutions for the hyperbolic SE kernel derivatives. Via multiple experiments, we demonstrated that hyperbolic pullback geodesics outperform both hyperbolic geodesics and Euclidean pullback geodesics.
It is worth noting that the benefits of pullback geodesics are most evident when the data exhibits smooth transitions. They become less effective when the data is inherently comprised of distinct clusters, as the pullback geodesics cannot cross high-energy regions, i.e., the data manifold boundaries. This paper focused on 2D and 3D hyperbolic latent spaces as they are easier to visualize and computationally less expensive than higher-dimensional latent spaces, while benefiting from the increased capacity of hyperbolic spaces to embed hierarchical data. Future work will investigate extensions to higher-dimensional hyperbolic latent spaces, which require to deal with more complex expressions for hyperbolic kernels. Additionally, the need for computing manual kernel derivatives is unsatisfactory. Future work will explore alternative autodifferentiation techniques built on KeOps [19] to overcome this practical issue. Finally, the approximation of the 2D hyperbolic SE kernel is computationally expensive. Performance could potentially be increased by exploring different sampling strategies, e.g., by sampling from a Rayleigh distribution rather than from a Gaussian.
@@ Appendix B Hyperbolic Kernel Derivatives
This section provides the complete derivations of the hyperbolic SE kernel derivatives that are necessary to compute the pullback metric tensor and to optimize geodesics on it. Eq. (16) defines the pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ based on the Jacobian mean $\bm{\mu}_{J}$ and covariance matrix $\bm{\Sigma}_{J}$ (15), which in turn require the kernel derivatives $\partial k(\bm{x}^{*},\bm{X})=\frac{\partial}{\partial\bm{x}^{*}}k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N}$ and $\partial^{2}k(\bm{x}^{*},\bm{x}^{*})=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}\in\mathbb{R}^{D_{x}\times D_{x}}$. Additionally, minimizing the curve energy (17), requires the derivative of the pullback metric tensor, i.e., the derivative of $\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J}$, which requires the kernel derivative $\dfrac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})$ and $\dfrac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})$. Specifically, we have
@@
where we have $\bm{\mu}_{J}\in\mathbb{R}^{D_{y}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}\in\mathbb{R}^{D_{y}\times D_{x}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N\times D_{x}}$, and ×$\times_{n}$ denotes the $n$-th mode tensor product [29]. Next, we provide the details of the derivation of the kernel derivatives for $\mathbb{H}^{2}_{\mathcal{L}}$ and $\mathbb{H}^{3}_{\mathcal{L}}$.
@@
We compute the 2D hyperbolic SE kernel $k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})$ via the Monte Carlo approximation (4). The pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ is computed using the kernel derivatives $\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{x}^{*}}$ and $\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$ (14). The only part of the kernel (4) that depends on the inputs $\bm{x},\bm{z}$ is the function $\Phi_{l}(\bm{x},\bm{z})$. Therefore, the derivatives of the 2D hyperbolic SE kernel (4) are entirely determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$, which are given by,
$\dfrac{\partial\Phi_{l}(\bm{x},\bm{z})}{\partial\bm{x}}=\left[\dfrac{\partial\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}}\right]\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})\,\quad\text{and}\quad\dfrac{\partial^{2}\Phi_{l}(\bm{x},\bm{z})}{\partial\bm{z}\partial\bm{x}}=\left[\dfrac{\partial\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}}\right]\left[\dfrac{\partial\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})}{\partial\bm{z}}\right]^{\mathsf{T}}\,.$ (35)
Eq (35) shows that these derivatives in turn are defined by the first and second derivatives of $\phi_{l}(\bm{x}_{\mathcal{P}})=e^{(1+2s_{l}i)\langle\bm{x}_{\mathcal{P}},\bm{b}_{l}\rangle}$, which are computed as
$\dfrac{\partial\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}}$ $=(1+2s_{l}i)\,\phi_{l}(\bm{x}_{\mathcal{P}})\left[\dfrac{\partial}{\partial\bm{x}}\bm{x}_{\mathcal{P}}\right]^{\mathsf{T}}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right],$ (36)
$\dfrac{\partial^{2}\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}^{2}}$ $=(1+2s_{l}i)\Bigg{(}\left[\dfrac{\partial}{\partial\bm{x}}\bm{x}_{\mathcal{P}}\right]^{\mathsf{T}}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right]\left[\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}})\right]^{\mathsf{T}}$ (37)
@@
$\left(\dfrac{\partial}{\partial\bm{x}}\bm{x}_{\mathcal{P}}\right)_{i,j}$ =$=\begin{cases}-\frac{x_{i+1}}{(1+x_{0})^{2}}&j=0\\ \quad\frac{1}{1+x_{0}}&j=i+1\\ \quad 0&\text{otherwise}\end{cases}\quad,\quad\left(\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\bm{x}_{\mathcal{P}}\right)_{i,j,k}=\begin{cases}\quad\frac{2x_{i+1}}{(1+x_{0})^{3}}&j=k=0\\ -\frac{1}{(1+x_{0})^{2}}&j=0\text{ and }k=i+1\\ -\frac{1}{(1+x_{0})^{2}}&k=0\text{ and }j=i+1\\ \quad 0&\text{otherwise}\end{cases}\,,$ (41)
@@
Minimizing the curve energy (17) to compute pullback geodesics additionally requires the derivative $\frac{\partial}{\partial\bm{x}^{*}}\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. This, in turn, requires the derivative of the Jacobian of the covariance matrix $\frac{\partial\bm{\Sigma}_{J}}{\partial\bm{x}^{*}}$, which depends on the kernel derivatives $\frac{\partial^{3}k(\bm{x},\bm{z})}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}|_{\bm{x}=\bm{z}=\bm{x}^{*}}$, and $\frac{\partial^{3}k(\bm{x},\bm{z})}{\partial\bm{z}^{2}\partial\bm{x}}|_{\bm{x}=\bm{z}=\bm{x}^{*}}$. Again, the kernel derivatives are determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$,
$\frac{\partial^{3}\Phi_{l}(\bm{x},\bm{z})}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}=\left[\dfrac{\partial^{2}\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}^{2}}\right]\times_{2}\left[\dfrac{\partial\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})}{\partial\bm{z}}\right]\quad\text{and}\quad\frac{\partial^{3}\Phi_{l}(\bm{x},\bm{z})}{\partial\bm{z}^{2}\partial\bm{x}}=\left[\dfrac{\partial^{2}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})}{\partial\bm{z}^{2}}\right]\times_{1}\left[\dfrac{\partial\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}}\right],$ (42)
where $\bm{A}\times_{i}\bm{B}$ refers to n-mode product [30] which multiplies the $i$-th dimension of $\bm{A}$ with the second dimension of $\bm{B}$. If $\bm{B}$ is a vector, $\bm{A}$ is unsqueezed at dimension $i$ before the multiplication. The matrices $\frac{\partial^{2}\phi_{l}(\bm{x}_{\mathcal{P}})}{\partial\bm{x}^{2}},\frac{\partial^{2}\phi_{l}(\bm{z}_{\mathcal{P}})}{\partial\bm{z}^{2}}\in\mathbb{R}^{3\times 3}$ in (42) are therefore interpreted as $\mathbb{R}^{3\times 1\times 3}$ and $\mathbb{R}^{1\times 3\times 3}$, respectively.
@@
This section provides the derivations of the 3D hyperbolic SE kernel derivatives. The 3D hyperbolic SE kernel is given by (3). In this section, we simplify its notation and write $k(\bm{x},\bm{z})=\frac{\rho}{s}e^{-\frac{\rho^{2}}{\nu}}$ by omitting the constant $\tau/C_{\infty}$, setting $\nu=2\kappa^{2}$, and introducing the helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$ and $s=\sqrt{u^{2}-1}$. Their derivatives are $\frac{\partial u}{\partial\bm{x}}=\bm{G}^{\mathcal{L}}\bm{z}$, $\frac{\text{d}\rho}{\text{d}u}=\frac{-1}{s}$, and $\frac{\text{d}s}{\text{d}u}=\frac{u}{s}$. Using these expressions, the first and second derivatives of the 3D hyperbolic SE kernel $k(\bm{x},\bm{y})=\frac{\rho}{s}\,e^{-\frac{\rho^{2}}{\nu}}$ are given as,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}\rho\right]\dfrac{1}{s}e^{-\frac{\rho^{2}}{\nu}}+\rho\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}s^{-1}\right]e^{-\frac{\rho^{2}}{\nu}}+\dfrac{\rho}{s}\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\nu}}\right],$
$=-\dfrac{1}{s^{2}}\bm{G}^{\mathcal{L}}\bm{z}e^{-\frac{\rho^{2}}{\nu}}-\dfrac{\rho u}{s^{3}}\bm{G}^{\mathcal{L}}\bm{z}e^{-\frac{\rho^{2}}{\nu}}+\dfrac{\rho}{s^{2}}\bm{G}^{\mathcal{L}}\bm{z}\dfrac{2\rho}{\nu}e^{-\frac{\rho^{2}}{\nu}},$
@@
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\bm{Gz}\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\nu}}+\bm{Gz}g(u)\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\nu}}\right]+g(u)e^{-\frac{\rho^{2}}{\nu}}\left[\dfrac{\partial}{\partial\bm{z}}\bm{Gz}\right],$
$=\left[\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\nu}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)\dfrac{-1}{s}\dfrac{-2\rho}{\nu}e^{-\frac{\rho^{2}}{\nu}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)e^{-\frac{\rho^{2}}{\nu}}\bm{G}^{\mathcal{L}},$
@@
$\dfrac{\partial^{2}}{\partial\bm{x}\partial\bm{x}}k(\bm{x},\bm{z})$ $=h(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{z}^{\mathsf{T}}\bm{G}^{\mathcal{L}}e^{-\frac{\rho^{2}}{\nu}}\,,$ (45)
with $g(u)\!=\!\left(\frac{2\rho^{2}}{\nu s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$ and $h(u)\!=\!\frac{\text{d}}{\text{d}u}g(u)+\frac{g(u)2\rho}{\nu s}$. The helper functions $g(u)$ and $h(u)$ are essential to understand why standard automatic differentiation tools fail to compute the derivatives (43)-(45). For equal inputs $\bm{x}\!=\!\bm{z}$, the inner product $u$ approaches $-1$, while the distance $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})$ and variable $s$ converge to 00. This limit is analytically well behaved for the helper functions, e.g., $\lim_{u\rightarrow-1^{-}}g(u)=\frac{2}{\nu}+\frac{1}{3}$. However, automatic differentiation tools fail to compute the kernel derivatives for equal inputs correctly because they do not cancel out 00-approaching terms. For example, we have for $g(u)$,
In this scenario, autodiff divides by 00, leading to undefined values (NaN). In other cases, e.g., when training a GPLVM, derivatives for equal kernel inputs $\bm{x}=\bm{z}$ may not be that relevant because different latent points rarely become that close. However, these derivatives are essential for computing the Jacobian covariance matrix $\bm{\Sigma}_{J}$ in (15). Although symbolic differentiation libraries could provide a solution, they tend to be too slow for practical purposes. To address this issue, we replace the general function definitions with their analytical limits, when $u$ gets close to $-1$. This is equivalent to the hyperbolic distance between the two kernel inputs falling below a predefined threshold value $\rho=\text{arccosh}(-u)=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})\leq 1\mathrm{e}{-4}$. The analytical limit expressions for the helper functions are given as
Finally, we need the third-order kernel derivatives to optimize pullback geodesics as in the 2D case. Denoting the second kernel derivative as $\bm{K}=\frac{\partial^{2}k(\bm{x},\bm{z})}{\partial\bm{z}\partial\bm{x}}\in\mathbb{R}^{4\times 4}$, the third-order derivatives is obtained by stacking the individual derivatives of each row and column of $\bm{K}$. The derivative of each row $\bm{K}_{i}$ w.r.t. the first input $\bm{x}$ and each column $\bm{K}_{j}$ w.r.t the second input $\bm{z}$ are given by,
@@
$\dfrac{\partial}{\partial\bm{x}}\bm{K}_{i}$ $=\left(q(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}z_{i}\,+h(u)\bm{G}^{\mathcal{L}}z_{i}\,+h(u)\bm{Gz}\bm{e}_{i}^{\mathsf{T}}\right)\Delta_{1i}e^{-\frac{\rho^{2}}{\nu}}\,,$ (53)
where $q(u)=\frac{\text{d}}{\text{d}u}h(u)+\frac{h(u)2\rho}{\nu s}$ and $\Delta_{ij}$ represents a modified Kronecker delta equal to $-1$ for equal indices and to $1$ otherwise. Stacking the row and column matrices $\bm{K}_{i}$ and $\bm{K}_{j}$ along the first and second dimension accordingly gives the third kernel derivatives $\frac{\partial\bm{K}}{\partial\bm{x}},\frac{\partial\bm{K}}{\partial\bm{z}}\in\mathbb{R}^{4\times 4\times 4}$.